# Polynomial-Time Near-Optimal Estimation over Certain Type-2 Convex Bodies

### v1

@@ 3 Near-optimal Estimation Algorithm
We make several comments about the algorithm. Note that, since we can evaluate the Minkowski gauge of $K$, we have a (weak) separation oracle for $K$ grotschel2012geometric, which in turn enables (weak) projection in polynomial-time (see Theorem 2.5.9 in dadush2012integer, see also lee2018efficient). The properties of this weak projection are summarized in Lemma A.1, where it is observed that it can output a vector that is in $K$ which is closer (up to an absolute constant factor) to any vector in the set $K$ than the original vector. When we write $\Pi_{K}$ we mean using this weak projection. In addition, for any $j$ the set $K_{(j)}$ also admits weak projection. This is because the Minkowski gauge of $K_{(j)}$ satisfies
@@
We will now explicitly write out the first couple of iterations of our algorithm, to ease the reader in the inductive logic. Every time, we will condition on the event $\|\tilde{\mu}_{i+1}-(\mu-\hat{\mu}_{i})/2\|_{2}\leq\tilde{d}_{i}/L$. In the end we will use the union bound using logic similar to Lemma B.5 of prasadan2024information to place ourselves on all of these events.
@@
Second, it is possible that, $\tilde{d}_{k}\leq(2r)\vee(C\sigma)$ and hence we have returned $\hat{\mu}_{k}$. First suppose $2r>C\sigma$. In that case, recall that we are assuming $\sigma\geq r/\sqrt{n}$ (otherwise, we simply output $Y$), implying by Lemma 2.10 that the minimax rate $\eta^{\star}(K)\gtrsim r$. On the other hand, since we are assuming $\|\hat{\mu}_{k}-\mu\|_{2}\lesssim\tilde{d}_{k}$, $\hat{\mu}_{k}$ is rate optimal. Next, suppose $C\sigma\geq 2r$. In that case since $\eta^{\star}(K)\gtrsim(C\sigma)$ (see prasadan2024some e.g.) it follows that $\hat{\mu}_{k}$ is rate optimal.
@@
Using steps similar to the proofs of Lemmas B.5, B.6 and B.7 of prasadan2024information along with Lemma 2.9, we can show that the accumulated probability bounds for $t\gtrsim\bigg[2\sqrt{2}T_{2}(K)\cdot\max_{k\in[M]}\kappa^{1/2}(K_{(k)})\vee\sqrt{4C^{2}\log_{c}(\max_{k\in[M]}\kappa^{1/2}(K_{(k)})\cdot\sqrt{2}T_{2}(K))}\bigg]\eta^{\star}(K)$, are:
@@
We remind the reader that in the inequality above, the bounds $\frac{1}{\tilde{d}_{i}^{3}/(\sigma^{3})}$ come from the potential failure of our oracle $\mathcal{O}_{K_{(j)}}(\cdot)$. We remark here that $r\leq\sqrt{n}\sigma$ (otherwise we simply output $Y$ which is optimal), implies that $\log R/\sigma$ is polynomial in $n$, ensuring that all iterations that we make in Section 3 to calculate the oracle $\mathcal{O}_{K_{(j)}}$ are polynomial in $n$ (recall Remark 2.6). Observe that the above bound continues to hold for values of $t\geq d$ as well (the RHS is actually simply 0 then, which is trivially upper bounded by the exponential). Note that for smaller values of $t$ we do not need any probability bounds since we have the trivial:
@@
where $\lesssim$ hides absolute constant factors and $\Upsilon(K)\geq 1$ depends on $T_{2}(K)$ and $\max_{k\in[M]}\kappa(K_{(k)})$ as described above. The final inequality holds true since, $\sigma^{2}\lesssim\eta^{\star 2}(K)$ (as shown in prasadan2024some e.g.).
@@ 3 Quadratic Form Maximization (QFM)
In this section, we review some of the results of bhattiprolu2021framework, and in addition, we establish some new results regarding QFM over convex bodies. Coupled with the results from Section 2 this demonstrates that our algorithm works unconditionally for plenty of type-2, well-balanced convex bodies.
@@
This is a diffuclt problem, but fortunately the paper bhattiprolu2021framework studied it in great detail and provided polynomial-time algorithms based on the ellipsoid algorithm for a plethora of origin symmetric type-2 convex bodies. In our case the matrix $X$ is symmetric psd. In bhattiprolu2021framework, the authors considered the same setting, and even the broader setting when $X$ is simply symmetric.
@@
In order to solve the two problems, bhattiprolu2021framework proceed to define the lower and upper covariance regions. The upper covariance region is defined as:
@@
bhattiprolu2021framework observe that
@@
where, as above, $K^{\circ}$ is the polar body of $K$. In that sense the upper and lower covariance sets are approximately convex. In their sections 3 and 4 the authors of bhattiprolu2021framework establish conditional polynomial-time algorithms. These algorithms, which are based on an (approximate) ellipsoid algorithms, require the existence of an approximate separation oracle for (at least) one of the two sets $\mathcal{U}(K),\mathcal{L}(K)$. The reader is kindly refered to Definitions 1.6 and 1.11 in bhattiprolu2021framework for what an approximate separation oracle means. The conditional [^10] Here the word conditional means that these algorithms work assuming separation oracles for one of the upper or lower covariance bodies. maximization polynomial-time algorithms guarantees are given in Propositions 4.2 and 4.4 of bhattiprolu2021framework. Importantly, we note that the algorithms in Propositions 4.2 and 4.4 not only can provide a near optimal value but also have rounding algorithms which produce near optimal vectors with high probability after polynomially many rounds. In addition to the approximate separation oracles for $\mathcal{U}(K)$ or $\mathcal{L}(K)$, these results require the set $K$ to be balanced in the sense that $rB_{2}\subseteq K\subseteq RB_{2}$ for some known constants $R,r\in\mathbb{R}_{+}$. The complexity of the algorithms is polynomial in $n,\log R/r$ and $\operatorname{bit}(X)$. Section 6 of bhattiprolu2021framework shows that the class of convex bodies for which we can evaluate an approximate separation oracle of the upper covariance region $\mathcal{U}(K)$ is closed under variety of operations including intersection, Minkowski addition, subspaces and quotients. This implies in particular that if one has a separation oracle for $\mathcal{U}(K)$ one is able to maximize over the set $K\cap cB_{2}$ for any $c\geq r$, since it is easy to see [^11] Indeed results in Section 7 of bhattiprolu2021framework show that any type-2 exactly 2-convex and sign-invariant norm like $\|x\|_{2}/c$ (the Minkowski gauge of $cB_{2}$), admits an approximate separation oracle of $\mathcal{U}(cB_{2})$, and hence by results of Section 6 the intersection set $K\cap cB_{2}$ also admits such an oracle). See also the proof of Theorem 3.3 where we explain why explicit separation oracle of the convex set $\mathcal{U}(cB_{2})$ exists. that the set $\mathcal{U}(cB_{2})$ admits a separation oracle. Next, in Section 7 bhattiprolu2021framework exhibit unconditional algorithms for a plethora of convex bodies $K$, where one has to have an oracle computing $\rho_{K}(x)$ for any $x\in\mathbb{R}^{n}$ (see Theorem 7.6 and Theorem Theorem 7.12). In other words, the authors are able to construct the approximate separation oracles for the upper and lower covariance regions of these sets. Specifically, the assumptions on the Minkowski gauges of the convex bodies in the latter theorems are as follows:
@@
Using the first relaxation given by bhattiprolu2021framework we can rewrite the problem as
@@
The proof is identical to Observation 4.1 of bhattiprolu2021framework. We have
@@
where the 1/2 is taken entry-wise on the non-negative vector $\operatorname{diag}(AWA^{\top})$. This is a convex function according to the proof of Theorem 7.6 of bhattiprolu2021framework. Specifically, the proof of Theorem 7.6 argues that the map $Y\mapsto\|\operatorname{diag}(Y)^{1/2}\|^{2}$ is convex for any psd matrix $Y\in\mathbb{R}^{m\times m}$. Since the map $Y\mapsto AYA^{\top}$ is affine: $A(\theta Y_{1}+(1-\theta)Y_{2})A^{\top}=\theta AY_{1}A^{\top}+(1-\theta)AY_{2}A^{\top}$, then it follows that the map $W\mapsto\|\operatorname{diag}(AWA^{\top})^{1/2}\|^{2}$ is also convex.
Using the proof of Theorem 7.6 (and Theorem 7.4) of bhattiprolu2021framework and the properties of $\|\cdot\|$, we can claim that
@@
We know by the properties of $W$ that $\langle X,W\rangle=\mathbb{E}\langle W^{1/2}g,XW^{1/2}g\rangle\geq(1-o(1))\operatorname{OPT}/(\beta\cdot T^{2}_{2}(K))$, where $\operatorname{OPT}$ is the optimal value of program (3.1). Next, following the idea of the rounding algorithm from Proposition 4.2 of bhattiprolu2021framework we generate Gaussian variables $g\in\mathbb{R}^{n}$ and compute, $G:=\langle W^{1/2}g,XW^{1/2}g\rangle-(1-o(1))\|AW^{1/2}g\|^{2}\operatorname{OPT}/((1+\varepsilon)\beta\cdot T^{2}_{2}(K))$, for some small $\epsilon>0$, until the number is positive for one of the vectors $g$. By Chebyshev’s inequality we can show that this will happen in polynomially many steps. Next, the approximately optimal solution is $W^{1/2}g/\|AW^{1/2}g\|$. We will now explicitly show the argument with Chebyshev’s inequality. Taking expectation we have
@@
where the penultimate inequality follows from Khintchine-Kahane’s theorem (see Theorem 2.1 in bhattiprolu2021framework e.g.), and $\delta:=1+\frac{1}{(1+\epsilon)\beta\cdot T_{2}^{2}(K))^{2}}$. Thus by Chebyshev’s inequality with probability at least $1-1/k^{2}$:
@@
Thus again by Corollary 4.12 part 1. we can claim, that we can find a matrix $W$ which is at most $2\beta\cdot T^{2}_{2}(K\cap cB_{2})\leq 4\beta\cdot\max(T^{2}_{2}(K),1)=4\beta\cdot T^{2}_{2}(K)$ approximate maximizer. Now to the rounding algorithm. The logic is the same as before: we know that $W$ that $\mathbb{E}\langle W^{1/2}g,XW^{1/2}g\rangle\geq(1-o(1))\operatorname{OPT}/(\beta\cdot T^{2}_{2}(K))$. We then generate polynomially many Gaussian variables $g\in\mathbb{R}^{n}$ and compute, $\langle W^{1/2}g,CW^{1/2}g\rangle-(1-o(1))(\rho_{K}(W^{1/2}g)\vee[\|W^{1/2}g\|_{2}c^{-1}])^{2}\operatorname{OPT}/((1+\varepsilon)\beta\cdot T^{2}_{2}(K))$. And the maximum $g$ gives a $W^{1/2}g/(\rho_{K}(W^{1/2}g)\vee[\|W^{1/2}g\|_{2}c^{-1}])$, which is our approximate solution. As before the probability can be boosted to 1q1-q with $c\log q^{-1}$ generations of Gaussians.
@@
Thus, we can approximately optimize the above problem by the results of Section 7 of bhattiprolu2021framework, see Theorem 7.12. Here $\kappa\lesssim T_{2}^{6}(K)\log^{4}(eT_{2}(K))$.
Using Proposition 4.13 of bhattiprolu2021framework we can now claim that we have a $\alpha\lesssim T_{2}^{8}(K)\log^{4}(eT_{2}(K))$-approximate separation oracle for the set $\mathcal{U}(K)$ where $K=\{x\in\mathbb{R}^{n}:\|Ax\|\leq 1\}$. Next, using Proposition 6.5 of bhattiprolu2021framework we know we can obtain a polynomial-time approximate quadratic maximization algorithm over the set $K\cap cB_{2}$ for any $r\leq c\leq R$. Not only that but we will also have rounding algorithms to obtain $\mathcal{O}(X)$ (we can simply replicate the rounding algorithms logic from the proof of Theorem 3.1 above which describes how to obtain rounding algorithms with the probability guarantees that we require). Here the constant $\kappa\lesssim 2\max(\alpha,1)\cdot\max(T^{2}_{2}(K),1)$.

### v2

@@ 3 Near-optimal Estimation Algorithm
We make several comments about the algorithm. Note that, since we can evaluate the Minkowski gauge of $K$, we have a (weak) separation oracle for $K$ Grötschel et al. (2012), which in turn enables (weak) projection in polynomial-time (see Theorem 2.5.9 in Dadush (2012), see also Lee et al. (2018)). The properties of this weak projection are summarized in Lemma A.1, where it is observed that it can output a vector that is in $K$ which is closer (up to an absolute constant factor) to any vector in the set $K$ than the original vector. When we write $\Pi_{K}$ we mean using this weak projection. In addition, for any $j$ the set $K_{(j)}$ also admits weak projection. This is because the Minkowski gauge of $K_{(j)}$ satisfies
@@
We will now explicitly write out the first couple of iterations of our algorithm, to ease the reader in the inductive logic. Every time, we will condition on the event $\|\tilde{\mu}_{i+1}-(\mu-\hat{\mu}_{i})/2\|_{2}\leq\tilde{d}_{i}/L$. In the end we will use the union bound using logic similar to Lemma B.5 of Prasadan and Neykov (2024) to place ourselves on all of these events.
@@
Second, it is possible that, $\tilde{d}_{k}\leq(2r)\vee(C\sigma)$ and hence we have returned $\hat{\mu}_{k}$. First suppose $2r>C\sigma$. In that case, recall that we are assuming $\sigma\geq r/\sqrt{n}$ (otherwise, we simply output $Y$), implying by Lemma 2.10 that the minimax rate $\eta^{\star}(K)\gtrsim r$. On the other hand, since we are assuming $\|\hat{\mu}_{k}-\mu\|_{2}\lesssim\tilde{d}_{k}$, $\hat{\mu}_{k}$ is rate optimal. Next, suppose $C\sigma\geq 2r$. In that case since $\eta^{\star}(K)\gtrsim(C\sigma)$ (see Prasadan and Neykov (2025b) e.g.) it follows that $\hat{\mu}_{k}$ is rate optimal.
@@
Using steps similar to the proofs of Lemmas B.5, B.6 and B.7 of Prasadan and Neykov (2024) along with Lemma 2.9, we can show that the accumulated probability bounds for $t\gtrsim\bigg[2\sqrt{2}T_{2}(K)\cdot\max_{k\in[M]}\kappa^{1/2}(K_{(k)})\vee\sqrt{4C^{2}\log_{c}(\max_{k\in[M]}\kappa^{1/2}(K_{(k)})\cdot\sqrt{2}T_{2}(K))}\bigg]\eta^{\star}(K)$, are:
@@
We remind the reader that in the inequality above, the bounds $\frac{1}{1+\tilde{d}_{i}^{3}/(\sigma^{3})}$ come from the potential failure of our oracle $\mathcal{O}_{K_{(j)}}(\cdot)$. We remark here that $r\leq\sqrt{n}\sigma$ (otherwise we simply output $Y$ which is optimal), implies that $\log R/\sigma$ is polynomial in $n$, ensuring that all iterations that we make in Section 3 to calculate the oracle $\mathcal{O}_{K_{(j)}}$ are polynomial in $n$ (recall Remark 2.6). Observe that the above bound continues to hold for values of $t\geq d$ as well (the RHS is actually simply 0 then, which is trivially upper bounded by the exponential). Note that for smaller values of $t$ we do not need any probability bounds since we have the trivial:
@@
where $\lesssim$ hides absolute constant factors and $\Upsilon(K)\geq 1$ depends on $T_{2}(K)$ and $\max_{k\in[M]}\kappa(K_{(k)})$ as described above. The final inequality holds true since, $\sigma^{2}\lesssim\eta^{\star 2}(K)$ (as shown in Prasadan and Neykov (2025b) e.g.).
@@ 3 Quadratic Form Maximization (QFM)
In this section, we review some of the results of Bhattiprolu et al. (2021), and in addition, we establish some new results regarding QFM over convex bodies. Coupled with the results from Section 2 this demonstrates that our algorithm works unconditionally for plenty of type-2, well-balanced convex bodies.
@@
This is a diffuclt problem, but fortunately the paper Bhattiprolu et al. (2021) studied it in great detail and provided polynomial-time algorithms based on the ellipsoid algorithm for a plethora of origin symmetric type-2 convex bodies. In our case the matrix $X$ is symmetric psd. In Bhattiprolu et al. (2021), the authors considered the same setting, and even the broader setting when $X$ is simply symmetric.
@@
In order to solve the two problems, Bhattiprolu et al. (2021) proceed to define the lower and upper covariance regions. The upper covariance region is defined as:
@@
Bhattiprolu et al. (2021) observe that
@@
where, as above, $K^{\circ}$ is the polar body of $K$. In that sense the upper and lower covariance sets are approximately convex. In their sections 3 and 4 the authors of Bhattiprolu et al. (2021) establish conditional polynomial-time algorithms. These algorithms, which are based on an (approximate) ellipsoid algorithms, require the existence of an approximate separation oracle for (at least) one of the two sets $\mathcal{U}(K),\mathcal{L}(K)$. The reader is kindly refered to Definitions 1.6 and 1.11 in Bhattiprolu et al. (2021) for what an approximate separation oracle means. The conditional [^10] Here the word conditional means that these algorithms work assuming separation oracles for one of the upper or lower covariance bodies. maximization polynomial-time algorithms guarantees are given in Propositions 4.2 and 4.4 of Bhattiprolu et al. (2021). Importantly, we note that the algorithms in Propositions 4.2 and 4.4 not only can provide a near optimal value but also have rounding algorithms which produce near optimal vectors with high probability after polynomially many rounds. In addition to the approximate separation oracles for $\mathcal{U}(K)$ or $\mathcal{L}(K)$, these results require the set $K$ to be balanced in the sense that $rB_{2}\subseteq K\subseteq RB_{2}$ for some known constants $R,r\in\mathbb{R}_{+}$. The complexity of the algorithms is polynomial in $n,\log R/r$ and $\operatorname{bit}(X)$. Section 6 of Bhattiprolu et al. (2021) shows that the class of convex bodies for which we can evaluate an approximate separation oracle of the upper covariance region $\mathcal{U}(K)$ is closed under variety of operations including intersection, Minkowski addition, subspaces and quotients. This implies in particular that if one has a separation oracle for $\mathcal{U}(K)$ one is able to maximize over the set $K\cap cB_{2}$ for any $c\geq r$, since it is easy to see [^11] Indeed results in Section 7 of Bhattiprolu et al. (2021) show that any type-2 exactly 2-convex and sign-invariant norm like $\|x\|_{2}/c$ (the Minkowski gauge of $cB_{2}$), admits an approximate separation oracle of $\mathcal{U}(cB_{2})$, and hence by results of Section 6 the intersection set $K\cap cB_{2}$ also admits such an oracle). See also the proof of Theorem 3.3 where we explain why explicit separation oracle of the convex set $\mathcal{U}(cB_{2})$ exists. that the set $\mathcal{U}(cB_{2})$ admits a separation oracle. Next, in Section 7 Bhattiprolu et al. (2021) exhibit unconditional algorithms for a plethora of convex bodies $K$, where one has to have an oracle computing $\rho_{K}(x)$ for any $x\in\mathbb{R}^{n}$ (see Theorem 7.6 and Theorem Theorem 7.12). In other words, the authors are able to construct the approximate separation oracles for the upper and lower covariance regions of these sets. Specifically, the assumptions on the Minkowski gauges of the convex bodies in the latter theorems are as follows:
@@
Using the first relaxation given by Bhattiprolu et al. (2021) we can rewrite the problem as
@@
The proof is identical to Observation 4.1 of Bhattiprolu et al. (2021). We have
@@
where the 1/2 is taken entry-wise on the non-negative vector $\operatorname{diag}(AWA^{\top})$. This is a convex function according to the proof of Theorem 7.6 of Bhattiprolu et al. (2021). Specifically, the proof of Theorem 7.6 argues that the map $Y\mapsto\|\operatorname{diag}(Y)^{1/2}\|^{2}$ is convex for any psd matrix $Y\in\mathbb{R}^{m\times m}$. Since the map $Y\mapsto AYA^{\top}$ is affine: $A(\theta Y_{1}+(1-\theta)Y_{2})A^{\top}=\theta AY_{1}A^{\top}+(1-\theta)AY_{2}A^{\top}$, then it follows that the map $W\mapsto\|\operatorname{diag}(AWA^{\top})^{1/2}\|^{2}$ is also convex.
Using the proof of Theorem 7.6 (and Theorem 7.4) of Bhattiprolu et al. (2021) and the properties of $\|\cdot\|$, we can claim that
@@
We know by the properties of $W$ that $\langle X,W\rangle=\mathbb{E}\langle W^{1/2}g,XW^{1/2}g\rangle\geq(1-o(1))\operatorname{OPT}/(\beta\cdot T^{2}_{2}(K))$, where $\operatorname{OPT}$ is the optimal value of program (3.1). Next, following the idea of the rounding algorithm from Proposition 4.2 of Bhattiprolu et al. (2021) we generate Gaussian variables $g\in\mathbb{R}^{n}$ and compute, $G:=\langle W^{1/2}g,XW^{1/2}g\rangle-(1-o(1))\|AW^{1/2}g\|^{2}\operatorname{OPT}/((1+\varepsilon)\beta\cdot T^{2}_{2}(K))$, for some small $\epsilon>0$, until the number is positive for one of the vectors $g$. By Chebyshev’s inequality we can show that this will happen in polynomially many steps. Next, the approximately optimal solution is $W^{1/2}g/\|AW^{1/2}g\|$. We will now explicitly show the argument with Chebyshev’s inequality. Taking expectation we have
@@
where the penultimate inequality follows from Khintchine-Kahane’s theorem (see Theorem 2.1 in Bhattiprolu et al. (2021) e.g.), and $\delta:=1+\frac{1}{(1+\epsilon)\beta\cdot T_{2}^{2}(K))^{2}}$. Thus by Chebyshev’s inequality with probability at least $1-1/k^{2}$:
@@
Thus again by Corollary 4.12 part 1. we can claim, that we can find a matrix $W$ which is at most $2\beta\cdot T^{2}_{2}(K\cap cB_{2})\leq 4\beta\cdot\max(T^{2}_{2}(K),1)=4\beta\cdot T^{2}_{2}(K)$ approximate maximizer. Now to the rounding algorithm. The logic is the same as before: we know that $W$ that $\mathbb{E}\langle W^{1/2}g,XW^{1/2}g\rangle\geq(1-o(1))\operatorname{OPT}/(\beta\cdot T^{2}_{2}(K))$. We then generate polynomially many Gaussian variables $g\in\mathbb{R}^{n}$ and compute, $\langle W^{1/2}g,CW^{1/2}g\rangle-(1-o(1))(\rho_{K}(W^{1/2}g)\vee[\|W^{1/2}g\|_{2}c^{-1}])^{2}\operatorname{OPT}/((1+\varepsilon)\beta\cdot T^{2}_{2}(K))$. And the maximum $g$ gives a $W^{1/2}g/(\rho_{K}(W^{1/2}g)\vee[\|W^{1/2}g\|_{2}c^{-1}])$, which is our approximate solution. As before the probability can be boosted to 1q1-q with $c\log q^{-1}$ generations of Gaussians. ∎
@@
Thus, we can approximately optimize the above problem by the results of Section 7 of Bhattiprolu et al. (2021), see Theorem 7.12. Here $\kappa\lesssim T_{2}^{6}(K)\log^{4}(eT_{2}(K))$.
Using Proposition 4.13 of Bhattiprolu et al. (2021) we can now claim that we have a $\alpha\lesssim T_{2}^{8}(K)\log^{4}(eT_{2}(K))$-approximate separation oracle for the set $\mathcal{U}(K)$ where $K=\{x\in\mathbb{R}^{n}:\|Ax\|\leq 1\}$. Next, using Proposition 6.5 of Bhattiprolu et al. (2021) we know we can obtain a polynomial-time approximate quadratic maximization algorithm over the set $K\cap cB_{2}$ for any $r\leq c\leq R$. Not only that but we will also have rounding algorithms to obtain $\mathcal{O}(X)$ (we can simply replicate the rounding algorithms logic from the proof of Theorem 3.1 above which describes how to obtain rounding algorithms with the probability guarantees that we require). Here the constant $\kappa\lesssim 2\max(\alpha,1)\cdot\max(T^{2}_{2}(K),1)$.